Week of 160509

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: David (databases), Krystof (computing), Maarten (SCOD + ALICE), Marian (monitoring + networks), Xavi (storage)
  • remote: David (FNAL), Di (TRIUMF), Dimitri (KIT), Ken (CMS), Kyle (OSG), Pepe (PIC), Rolf (IN2P3), Salvatore (CNAF), Tiju (RAL), Ulf (NDGF), Vincenzo (EGI), Zoltan (LHCb)

Experiments round table:

  • CMS reports (raw view) -
    • Really no issues to speak of that I have noticed.
    • Usual activities continue, including MC production and T0 processing.

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Activity
      • Pre-staging of 2012 data for incremental stripping is ongoing
      • Ongoing Data striping on 2012 data
      • two new validation productions (FULL & TURBO stream) on Collision16 data
    • Site Issues
      • T0: EOS problem during the long weekend is fixed(cannot upload and download data)
        • Xavi: we only saw clients disconnecting; there might have been some network issue
      • T1:
        • CNAF - SRM was returning 0 free space on the tape space token (fixed).
        • GRIDKA - 24 hours pre-staging lost, because all tape recalls where suspended (GGUS:121280)

Sites / Services round table:

  • ASGC:
  • BNL:
  • CNAF: ntr
  • FNAL: ntr
  • GridPP:
  • IN2P3:
    • There are several files of CMS with size 0. This is a remainder of an old bug in a site's script (corrected since several months) where write errors on tape were silently ignored and the files cataloged as saved (GGUS:121230)
  • JINR:
  • KISTI:
  • KIT:
    • we are in downtime until Wed evening for maintenance of the cooling infrastructures
  • NDGF: ntr
  • NL-T1:
  • NRC-KI:
  • OSG: ntr
  • PIC:
    • the T10KD drives and tapes are back in use for ATLAS
      • a few months ago they were excluded because of a problem that is understood now
      • a few tapes written by one drive were found to have their files misaligned
      • a Service Incident Report will soon be provided for the archive
  • RAL: ntr
  • TRIUMF: ntr

  • CERN computing services:
    • An intervention on Argus (upgrade to 1.7) planned for tomorrow will be delayed because of a code issue discovered by CMS.
  • CERN storage services:
    • EOS-ALICE incident this morning:
      • the head node fail-over machine was still running with only half the memory
      • the 2nd big-memory machine would finally take over that role
      • the namespace needed to be compacted first to allow for a fast transition
      • unfortunately it ran into an inconsistency and aborted
      • the inconsistency then was repaired manually
      • a reboot of the namespace then took 2h as usual (will be much less in future versions)
      • to avoid further downtime, the switch of the machines was postponed
        • it has been agreed for Tue morning
  • CERN databases:
    • Due to a disruptive intervention on network switches described at OTG0030304, the following databases will be unavailable on Wednesday, 11th May from 8:00 to 10:00
      • ATLAS integration databases INTR and INT8R
      • Castor integration databases
      • CMS integration databases INT2R and INT9R
      • LHCb integration database INT12R and general purpose physics integration database INTDB11 (aka TEST2)
      • WLCG integration databases INT6R and INT11R
  • GGUS:
  • Monitoring:
    • the SAM tests for ATLAS will be moved to the new ETF system on Tue
      • for the other experiments this move was already done 2 weeks ago
  • MW Officer:
  • Networks:
  • Security:

AOB:

  • ATTENTION: next meeting on Tuesday May 17 !
Edit | Attach | Watch | Print version | History: r10 < r9 < r8 < r7 < r6 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r10 - 2016-05-09 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback